OpenStudy (anonymous):

Find the 6th term of a geometric sequence t3 = 444 and t7 = 7104. So I have to find r. But is this right: 7104 = 444r^4 r^4 = 16 r = 2 Or should it be r^3? I'm never sure if the power is = the number of terms missing or something completely unrelated.

ganeshie8 (ganeshie8):

you got the value of r correctly. but the method you used seems wrong

ganeshie8 (ganeshie8):

im just lil curious... can you pls explain how you put this :- 7104 = 444r^4

OpenStudy (anonymous):

I have no idea what formula I used I just followed the example of another problem (from the pic)

OpenStudy (anonymous):

and since they didn't explain how they got to \[r^{2}\] I just assumed it was the number of terms or something. But I'm not sure

OpenStudy (phi):

t3 = 444 and t7 = 7104 you could figure it like this: multiply a term by r to get the next term: t4= t3*r t5= t4*4 = (t3*r)*r= t3*r^2 t6= t5*r= t3*r^3 t7 = t6*r= t3*r^4 to get from t3 to t4 you multiply by r once to get from t3 to t7 you multiply by r four times, r^4

OpenStudy (anonymous):

let a be the first term ,r the common ratio. tn=ar^(n-1) t3=ar^2=444 t7=ar^6=7104 divide \[\frac{ ar ^{6} }{ar ^{2} }=\frac{ 7104 }{444 }=16\] r^4=2^4 r=2 a*2^2=444 a=444/4 =111 t6=ar^5=111*2^5=111*32=3552 or t6*r=t7 t6=t7/r=7104/2=3552

OpenStudy (anonymous):

phi, that makes it very clear. Thanks a lot!